OpenStudy (anonymous):

This is a related rates problem [calculus 1] A fire has started in a dry, open field and spreads in the form of a circle. The radius of the circle increases at the rate of 6ft/min. Find the rate at which the fire is increasing when the radius is 150 feet.

OpenStudy (anonymous):

whats the formula for the area of a circle

OpenStudy (anonymous):

no pain, no gain

OpenStudy (anonymous):

a=pie r^2

OpenStudy (radar):

\[A=pir ^{2}\] d/dt[A]=d/dt[pi r^2]

OpenStudy (radar):

dA/dt=2 pi r(dr/dt)

OpenStudy (radar):

Do you follow so far?

OpenStudy (radar):

dA/dt = (2) (Pi) (150) (6) Your answer will be in square ft per minute.

OpenStudy (anonymous):

yes

OpenStudy (radar):

5,655 sq ft/min

OpenStudy (radar):

Seem like it is spreading pretty fast, maybe someone else can verify this answer.